### v2

Abstract
Pneumonia, a widespread respiratory infection, remains a leading cause of morbidity and mortality globally, especially in vulnerable populations. Chest X-rays are a primary tool in detecting pneumonia; however, differences in imaging conditions and subtle visual indicators can make consistent interpretation challenging. Automated tools can complement traditional methods by providing additional insights to enhance diagnostic reliability and support clinical decision-making. In this study, we propose a novel approach that leverages deep learning with integrated self-attention mechanisms to enhance pneumonia detection in chest X-rays. Our method begins with lung segmentation using a TransUNet model that integrates our specialized transformer module, which has fewer parameters compared to common transformers while maintaining performance. This model is trained on the ”Chest Xray Masks and Labels” dataset and then applied to the ”Kermany” and ”Cohen” datasets to generate lung regions, enhancing subsequent classification tasks. By isolating the lung areas, the segmentation step provides focused input for the classification stage, enabling the model to concentrate on relevant regions. For classification, we utilize pre-trained ResNet models (ResNet-50 and ResNet-101) specifically to extract multi-scale feature maps, a core aspect of our approach. These feature maps are then processed through our modified transformer module to enhance pneumonia detection. By employing our specialized transformer, we achieve superior results with significantly fewer parameters compared to common transformer models. Our approach achieves high accuracy rates of 92.79% on the ”Kermany” dataset and 95.11% on the ”Cohen” dataset, ensuring robust and efficient performance suitable for resource-constrained environments. https://github.com/amirrezafateh/Multi-Scale-Transformer-Pneumonia
Keywords:
Transformer, Multi Scale, Pneumonia, Classification
1 Introduction
A serious respiratory condition known as pneumonia results in inflammation of one or both lungs, which can cause fever, coughing, and breathing difficulties. Since it causes about 15% of mortality in children under five, it is especially risky for young children unicef_pneumonia . This illness is widespread in developing and impoverished nations, where access to healthcare is restricted, and situations like pollution, overcrowding, and subpar living conditions worsen the situation who_pneumonia_in_children .
In order to effectively treat pneumonia and improve patient outcomes, early diagnosis is essential. Still, identifying pneumonia can be challenging. It can be easily confused with other lung diseases, and making consistent and accurate diagnosis difficult health_poverty_action_poverty_and_health . Conventional imaging methods are frequently employed, including computed tomography (CT), magnetic resonance imaging (MRI), and chest radiographs. But, chest X-rays are favored for being non-invasive and cost-effective radiologyinfo_chest_xray .
The variability in interpreting chest X-ray images underscores the need for consistent, automated diagnostic tools. The development of deep learning technology offers a promising approach to addressing the challenges in pneumonia diagnosis gayathri2022computer ; lin2023enhancing ; ortiz2022automatic . As a branch of artificial intelligence (AI), deep learning entails training neural networks on extensive datasets to identify patterns and generate predictions. In medical imaging, deep learning algorithms can scrutinize chest X-rays with remarkable precision, often exceeding human radiologists in terms of consistency and speed. Convolutional Neural Networks (CNNs), a specialized deep learning model designed for image analysis, have been developed to identify pneumonia from chest X-rays askari2024enhancing . These models are trained to recognize patterns and subtle features associated with pneumonia, contributing to consistent and reliable diagnostic performance that complements the expertise of radiologists krizhevsky2012imagenet ; parhami2023comparison . However, traditional CNN-based models still cannot achieve impressive accuracy. Recent advancements include the use of ensemble learning techniques, which combine predictions from multiple models to improve overall accuracy, and the integration of attention mechanisms, which help models focus on the most relevant parts of the image fateh2024advancing . Transformers have demonstrated significant potential in visual tasks. They excel at modeling long-range dependencies and identifying intricate patterns in medical images fateh2024msdnet ; mabrouk2022pneumonia .
Lung segmentation is frequently used in chest X-ray images to improve the accuracy of pneumonia detection. This technique isolates the lung regions, allowing segmentation algorithms to minimize noise and concentrate the analysis on the pertinent areas. The U-Net architecture, known for its encoder-decoder structure and skip connections, has been particularly effective in medical image segmentation, including lung segmentation in chest X-rays. Accurate segmentation allows subsequent classification models to better identify pneumonia, leading to improved diagnostic performance rajpurkar2017chexnet . Furthermore, advancements in segmentation techniques, such as integrating attention mechanisms and transformers, have improved segmentation accuracy. These improvements enable models to concentrate on the most critical parts of the lung, capturing intricate details and ensuring that even the subtle signs of pneumonia are detected zhang2023cams ; fateh2023enhancing . Effective lung segmentation thus provides a robust foundation for classification algorithms, enabling them to achieve higher accuracy and reliability in diagnosing pneumonia.
A significant challenge in training these deep learning models is the requirement for large, high-quality labeled datasets, which are often hard to acquire in the medical field. Transfer learning tackles this problem by utilizing models pre-trained on extensive datasets like ImageNet and then fine-tuning them for particular medical tasks using smaller datasets. This approach has significantly improved the accuracy of medical image classifications chen2017rethinking . Even with transfer learning, models trained in one domain often fail to perform well in significantly different domains, such as medical applications. This highlights the need for more advanced and higher-level features, which is why transformers are increasingly used rezvani2024fusionlungnet . Despite the benefits of using transformers and their achievements in segmentation and classification tasks in CXR images, they bring two new challenges: models are more complex and have many parameters. These complex models used more hardware resources and took longer to train.
We propose an innovation approach that leverages deep learning with integrated self-attention mechanisms to overcome the challenges of traditional transformers while maintaining lower model complexity. Our method begins with lung segmentation using a TransUNet model, which integrates transformer-based attention mechanisms into the U-Net architecture. The TransUNet model is trained on the ”Chest Xray Masks and Labels” dataset candemir2013lung ; jaeger2013automatic to accurately segment lung regions in the images. Once trained, this pre-trained model is used with frozen weights to predict lung masks for our target datasets, ”Kermany” kermany2018labeled and ”Cohen” cohen2020covid . This segmentation step isolates the lung regions, enhancing the subsequent classification task.
For classification, we utilize pre-trained ResNet models, specifically ResNet-50 and ResNet-101, as the foundation for feature extraction. By drawing multi-scale feature maps from various stages of the ResNet models, we can leverage multiple feature spaces, which enhances the accuracy of our detection. This is achieved through a customized transformer module that employs a cross-attention mechanism, allowing us to make decisions based on more than one feature space. This transformer has been optimized to minimize the number of parameters while preserving performance. By concentrating on the relevant lung regions and integrating multi-scale information, our approach strives to achieve high diagnostic accuracy for pneumonia detection. This architecture reduces the computational load and ensures robust and reliable performance, making it suitable for deployment in resource-limited settings.
Our proposed method offers the following key contributions:
  • Development of a novel transformer structure that significantly reduces complexity compared to traditional transformer-based models while maintaining high performance.
  • Introduction of a novel TransUNet architecture for the segmentation task.
  • Incorporation of multi-scale feature extraction, enabling enhanced performance through the utilization of multiple feature spaces.
  • Reduced number of parameters compared to other state-of-the-art models.
  • Achieving high accuracy rates of 92.79% on the ”Kermany” dataset and 95.11% on the ”Cohen” dataset.
2 Related Work
In recent years, the focus of research on diagnosing and categorizing lung diseases, including pneumonia, through medical imaging has intensified, driven by advances in machine learning and deep learning technologies. Precisely segmenting lung areas in chest X-ray (CXR) images is essential for reliable disease identification and thorough analysis. This section examines deep learning techniques for segmenting and diagnosing lung diseases in chest X-ray (CXR) images. For the segmentation task, we focus on the U-Net architecture and its variations, including attention mechanisms and transformer blocks, which have significantly advanced lung disease segmentation. In the classification task, we categorize approaches into basic deep learning models, transfer learning, fine-tuning, and custom models, emphasizing how these advanced techniques have progressively improved diagnostic outcomes.
2.1 Segmentation
2.1.1 U-Net for CXR Segmentation
The U-Net architecture, with its encoder-decoder structure and skip connections, has occurred as a leading method for CXR segmentation. This setup, which captures high-level semantic information and low-level details, is crucial for accurately outlining lung boundaries. Studies have consistently shown U-Net’s effectiveness in segmenting lung regions with high accuracy, a factor that significantly comforts the potential of this technology in improving diagnostic outcomes fateh2023persian ; fateh2022providing . U-Net, introduced by Ronneberger et al., has become a fundamental tool in medical image segmentation ronneberger2015u . Islam et al. islam2018towards showcased U-Net’s ability in accurately tracing lung boundaries, which has enhanced diagnostic precision. Additionally, Liu et al. liu2022automatic employed a pre-trained EfficientNet-B4 and developed an enhanced version of U-Net for identifying and segmenting lung regions.
2.1.2 U-Net Enhancements with Transformers
Recent research has significantly advanced lung segmentation by enhancing the U-Net architecture with attention mechanisms. Oktay et al. oktay2018attention introduced mechanisms that enable the model to concentrate on the most crucial areas within chest X-rays using Attention Gates (AGs). This innovation enhances segmentation accuracy and sensitivity to disease characteristics. Additionally, research by Wu et al. wu2023transformer , Gu et al. gu2018recent , and Liu et al. liu2024sta has shown that incorporating attention mechanisms into the U-Net framework significantly improves lung segmentation performance, underscoring the effectiveness of this method in precisely outlining lung boundaries and enhancing diagnostic outcomes.
Khaniki et al. khaniki2024novel enhanced U-Net by incorporating a Convolutional Block Attention Module (CBAM), which integrates channel, spatial, and pixel attention to boost segmentation accuracy. Azad et al. and Chen et al. extended the U-Net framework with transformers, demonstrating significant improvements in capturing intricate details and achieving top-tier results in lung segmentation tasks azad2019bi ; chen2021transunet .
The incorporation of transformer modules has marked a landmark in lung segmentation research. Transformer architectures, known for capturing long-range dependencies and contextual information from text, have been successfully integrated into U-Net variants, leading to notable improvements in segmentation accuracy. For instance, Chen et al. chen2023cotrfuse created a hybrid CNN-Transformer model for medical image segmentation, merging the strengths of CNNs and transformers to enhance accuracy and robustness in lung tissue segmentation. Similarly, Valanarasu et al. valanarasu2021medical developed hybrid models combining U-Net with transformer modules, effectively utilizing CNNs and transformers to capture complex anatomical details and spatial relationships in chest X-ray images.
2.2 Classification
2.2.1 Classical Approaches for CXR Classification
Early methods for classifying chest X-ray (CXR) images primarily depended on traditional machine learning techniques, employing classifiers such as Support Vector Machines (SVM), K-nearest Neighbors (k-NN), and Random Forests. For example, Stokes et al. used logistic regression, decision trees, and SVM to categorize patients’ clinical data into bronchitis or pneumonia, with decision trees yielding the highest recall value of 80% and an AUC of 93% stokes2021machine . Similarly, Qi et al. applied logistic regression and random forest to CT images, achieving AUCs of 0.97 and 0.92, respectively qi2020machine . Chandra et al. used a multi-layer perceptron (MLP) to segment lung regions from CXR images, reaching an accuracy of 95.39% chandra2020pneumonia . However, these methods, which heavily relied on symptomatic data, had limited accuracy and were evaluated on small datasets wang2024prediction ; berg2023employing ; fateh2021multilingual .
2.2.2 Deep Learning Models
The beginning of deep learning, especially Convolutional Neural Networks (CNNs), has significantly transformed medical image analysis by providing superior accuracy and robustness. For instance, Stephen et al. designed a custom CNN model from scratch, achieving a training accuracy of 95.31% and a validation accuracy of 93.73% stephen2019efficient . Similarly, Sharma et al. created a straightforward CNN architecture that reached a 90.68% accuracy rate on the ”Kermany” dataset using data augmentation sharma2020feature ; kermany2018labeled . However, relying solely on data augmentation does not introduce substantially new information, restricting the model’s ability to learn a wide range of complex patterns from the training data.
2.2.3 Transfer Learning
Pre-trained CNNs have become the standard for image classification tasks, including CXR analysis. These models leverage large datasets and transfer learning to enhance performance on specific medical imaging tasks. Transfer learning, where pre-trained models are adapted and refined for new, specific tasks, has achieved significant results. For instance, Rajpurkar et al. utilized DenseNet-121 on the ChestX-ray8 dataset, comprising 112,150 frontal CXR images, achieving an F1-score of 76.8%. This study highlighted the potential of transfer learning in medical image classification rajpurkar2017chexnet . Choudhary et al. used customized VGG19 and transfer learning to classify Optical Coherence Tomography (OCT), achieving 99.17% accuracy choudhary2023deep . Rahman et al. rahman2020transfer utilized transfer learning techniques exclusively, employing various CNN models pre-trained on ImageNet data to classify pneumonia.
2.2.4 Ensemble Approaches
Ensemble learning, which combines the outputs of multiple CNN models, has shown considerable promise. For instance, Ukwuoma et al. ukwuoma2023hybrid proposed two ensemble methods: ensemble group A (DenseNet201, VGG16, and GoogleNet) and ensemble group B (DenseNet201, InceptionResNetV2, and Xception). These models, followed by a self-attention layer and a multi-layer perceptron (MLP) for disease identification, achieved 97.22% accuracy for binary classification, and 97.2% and 96.4% for multi-class classification, respectively. Kundu et al. kundu2021pneumonia combined GoogLeNet, ResNet-18, and DenseNet-121 in an ensemble model, reaching an accuracy of 86.85%, albeit with high computational costs. Jaiswal et al. jaiswal2019identifying used a mask region-based CNN for pneumonia detection through segmentation, employing an ensemble of ResNet-50 and ResNet-101 for image thresholding. In the RSNA Pneumonia Detection Challenge, Gabruseva et al. gabruseva2020deep presented a deep learning framework using the single-shot detector RetinaNet with Se-ResNext101 encoders, achieving a mean average precision (mAP) of 0.26 through snapshot ensembling.
2.2.5 Transformers
Recent advancements in medical image classification have harnessed transformer architectures alongside deep learning, yielding impressive outcomes. Wang et al. wang2021transpath unveiled TransPath, a hybrid model merging CNN and transformer architectures, highlighting the potential of such integrations. They proved the efficacy of self-supervised pretraining on extensive datasets like TCGA and PAIP, followed by fine-tuning on specific medical image datasets, resulting in solid performance: 89.68% accuracy on MHIST, 95.85% on NCT-CRC-HE, and 89.91% on PatchCamelyon. Transformer-based models have garnered attention for their capacity to capture long-range dependencies in images. Wu et al. wu2022swin introduced a Swin Transformer-based model for pulmonary nodule classification, successfully adapting the architecture to the smaller scale of medical image datasets and achieving significant results. Dai et al. dai2021transmed investigated the use of transformers for multimodal medical image classification with their TransMed model, reaching 88.9% accuracy on the PGT dataset and 85% on the MRNet dataset. Leamons et al. leamons2022vision concentrated on breast cancer detection, comparing CNNs, RNNs, and Visual Transformers (VTs), and found that the VT model excelled with a 93% accuracy. Jang et al. jang2022m3t proposed M3T, a 3D medical image classifier that combines transformers with 2D and 3D CNNs, showcasing its effectiveness across various medical datasets.
3 Proposed method
Figure 1: The overview of proposed method
3.1 Overview
In this study, we propose a novel approach for segmentation and classification of Pneumonia Chest X-ray images by leveraging the power of deep learning and transformer-based attention mechanisms. Our method utilizes pre-trained ResNet models, specifically ResNet-50 and ResNet-101, as the backbone for feature extraction. These models are well-known for their ability to capture intricate patterns and features in images due to their deep architecture and residual connections.
Our approach begins with a segmentation step where we employ a TransUNet model, which integrates transformer-based attention mechanisms into the popular U-Net architecture. This model is trained on ”Chest Xray Masks and Labels” dataset candemir2013lung ; jaeger2013automatic to accurately segment lung regions in the images. By predicting masks for ”Cohen” dataset cohen2020covid using this pre-trained TransUNet, we can isolate the regions of interest, enhancing the subsequent classification task.
The segmentation step provides us with precise lung masks, ensuring that our classification model focuses on the relevant areas of the X-ray images. This preprocessing step is crucial for improving the overall accuracy of the system by reducing background noise and irrelevant features.
Our classification approach extracts multi-scale feature maps from three key stages of the ResNet models: the outputs of Block 2, Block 3, and Block 4. These stages provide a rich set of features at different scales, which are crucial for accurately identifying Pneumonia in chest X-rays. The extracted feature maps are then processed through a specialized transformer module that employs a cross-attention mechanism. This transformer enhances the feature representation by allowing the network to focus on the most relevant parts of the image.
After the cross-attention processing, the feature maps are concatenated to form a comprehensive representation of the input image. This combined feature map is subsequently fed into fully connected layers to perform the final classification. The overall architecture is designed to effectively integrate multi-scale information and attention mechanisms, thereby improving the classification accuracy. An overview of the proposed method is illustrated in Figure 1.
3.2 Segmentation task
In our proposed method, the segmentation task is pivotal for isolating lung regions in chest X-ray images, thereby enhancing the accuracy of pneumonia classification. For this purpose, we have designed a TransUNet model, which uniquely combines the strengths of the U-Net architecture with advanced techniques.
3.2.1 TransUNet Architecture
The TransUNet architecture can be divided into three main components: the encoder, the bottleneck, and the decoder. The encoder consists of a series of convolutional layers designed to capture hierarchical features from the input image. Each stage of the encoder includes a double convolution block, which performs two consecutive convolutions followed by batch normalization and ReLU activation. This setup helps in learning complex features at multiple levels. The encoder progressively reduces the spatial dimensions while increasing the depth of the feature maps through max-pooling operations.
At the bottleneck stage, the most abstract features of the input image are captured. This layer consists of a double convolution block. The bottleneck also incorporates an embedding layer and a positional encoding mechanism, which prepare the feature maps for the subsequent transformer module. The detailed structure and function of the transformer module will be discussed later in the classification subsection.
The decoder aims to reconstruct the segmented output by progressively upsampling the feature maps and concatenating them with corresponding encoder layers. Each upsampling step is followed by a double convolution block, which refines the features and reduces the number of channels. This structure helps in restoring the spatial resolution of the feature maps while retaining the detailed information captured by the encoder. The final layer of the decoder is a convolutional layer with a single output channel, which predicts the segmentation mask for the lung regions.
3.2.2 Training the TransUNet Model
Initially, we resize all images to 512x512 pixels. The TransUNet model is trained on the ”Chest Xray Masks and Labels” dataset, which provides paired X-ray images and corresponding lung masks. By training on this dataset, the TransUNet model learns to accurately segment lung regions in chest X-rays, ensuring that the subsequent classification step focuses on the relevant areas, thereby improving the overall accuracy of the system.
By effectively combining the robust feature extraction capabilities of the U-Net architecture with advanced processing techniques, the TransUNet model provides a powerful solution for the segmentation task in our proposed method.
3.2.3 Applying the Trained TransUNet to the Cohen Dataset
After successfully training the TransUNet model on the ”Chest Xray Masks and Labels” dataset, we utilize this pre-trained segmentation model to predict lung masks for the ”Cohen” dataset. This step is crucial for enhancing the accuracy of the subsequent classification task by focusing on the lung regions within the X-ray images.
The ”Cohen” dataset, which contains chest X-ray images, requires preprocessing to ensure that our classification model focuses on the most relevant regions. To achieve this, we apply the trained TransUNet model to segment the lung areas from these images. The ”Cohen” dataset is first preprocessed to match the input requirements of the TransUNet model. This involves standardizing the image dimensions and normalizing the pixel values to ensure consistency with the training data used for the TransUNet model.
Using the pre-trained TransUNet model, we generate lung masks for each X-ray image in the ”Cohen” dataset. The segmentation model outputs binary masks that highlight the lung regions while suppressing the background. By utilizing the pre-trained TransUNet model to segment the lung regions in the ”Cohen” dataset, we effectively preprocess the data to improve the performance of our classification model. This segmentation step filters out noise and irrelevant features, allowing the classifier to concentrate on the lung areas, thereby enhancing the overall accuracy and robustness of our proposed method.
3.3 classification task
Following the segmentation of lung regions using the TransUNet model, the next step in our proposed method is the classification task. This task involves accurately identifying the presence of pneumonia or covid19 in the preprocessed chest X-ray images. By focusing on the lung regions isolated during the segmentation phase, we enhance the classification model’s ability to detect relevant features indicative of pneumonia or covid19, thereby improving diagnostic accuracy.
3.3.1 Backbone
The backbone of our proposed method utilizes a pre-trained ResNet model, specifically ResNet-50 or ResNet-101, to extract multi-scale feature maps from the input chest X-ray images. Initially, we resize all images to 512x512 pixels. We focus on the outputs from Blocks 2, 3, and 4 of the ResNet, denoted as $B^{2}$, $B^{3}$, and $B^{4}$ respectively. Each of these blocks provides feature maps of size $(c,h,w)$, where $h=w=64$ The channels $c$ for these blocks are 512, 1024, and 2048 respectively.
To handle the complexity and standardize the feature maps for subsequent processing, we apply 1x1 convolution operations to reduce the number of channels. Specifically, for the output of Block 4 ($B^{4}$), as shown in Equation 1, we reduce the channels to 64 using a 1x1 convolution.
$B^{\prime 4}=C_{1\times 1}(B^{4})$ (1)
where $C_{1\times 1}$ denotes the 1x1 convolution operation.
For the outputs of Blocks 2 and 3 ($B^{2}$ and $B^{3}$), as shown in Equation 2 and Equation 3, we use separate 1x1 convolutions to reduce the number of channels for each to 32.
$B^{\prime 2}=C_{1\times 1}(B^{2})$ (2)
$B^{\prime 3}=C_{1\times 1}(B^{3})$ (3)
where $C_{1\times 1}$ in each equation indicates a reduction in the number of channels to 32.
After reducing the channels, we concatenate the feature maps from Block 2 and Block 3 to form a merged feature map (Equation 4).
$B^{merged}=Cat(B^{\prime 2},B^{\prime 3})$ (4)
where $Cat$ denotes the concatenation operation.
Thus, we have two main feature maps with size of (64,64,64) for further processing:
  • $B^{\prime 4}$
  • $B^{merged}$
Figure 2: The overview of transformer
3.3.2 Transformer
In our proposed method, a transformer is employed to enhance the feature representation obtained from the ResNet backbone, and a similar transformer architecture is used within the TransUNet model. For the classification task, we leverage this transformer to refine the multi-scale feature maps extracted from the ResNet backbone. We begin with two feature maps of size (64,64,64) derived from the ResNet backbone: $B^{\prime 4}$ and $B^{merged}$. Each of these feature maps is fed into a separate transformer, although the structure of the transformers is identical. For simplicity, we will describe the process for $B^{\prime 4}$.
Global Average Pooling (GAP)
We apply global average pooling to $B^{\prime 4}$ to create a feature vector $V^{4}$ (Equation 5).
$V^{4}=GAP(B^{\prime 4})$ (5)
This vector serves as the query for the transformer.
Reshaping for Key and Value
The feature map $B^{\prime 4}$ is reshaped to form the key and value inputs for the transformer. Specifically, as shown in equation 6, $B^{\prime 4}$ is reshaped from $(c,h,w)$ to $(c,h\times w)$.
$B_{flat}^{{}^{\prime}4}=Reshape(B^{\prime 4})$ (6)
Attention Mechanism
Query($Q$): The query is obtained from the feature vector $V^{4}$ created by global average pooling.
Key($K$) and Value($V$): Both the key and value are derived from the reshaped feature map $B_{flat}^{{}^{\prime}4}$
Scaled Dot-Product Attention
The attention scores are computed as the dot product of the query and key, followed by a softmax operation to obtain the attention weights (Equation 7).
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{Q\cdot K^{T}}{\sqrt{d_{k}}}\right)\cdot V$ (7)
where $d_{k}$ is the dimensionality of the key, and $Q\cdot K^{T}$ represents the dot product of the query and the transposed key. The result is a weighted sum of the value vectors, producing an output feature vector $F^{4}$ of size equal to the channel dimension.
3.3.3 Output Feature Vector
The output of the transformer for $B^{\prime 4}$ is a feature vector $F^{4}$ of size equal to the number of channels (64 in this case).
The same process is applied to $B^{merged}$, resulting in another feature vector $F^{merged}$ of size 64. The entire process is visually represented in Figure 2, providing a detailed overview of the transformer’s operation on the feature maps.
3.3.4 Find correct class
After processing the feature maps $F^{4}$ and $F^{merged}$ through transformers, we concatenate these outputs to form a unified feature representation (Equation 8).
$F^{concat}=Cat(F^{4},f^{merged})$ (8)
The concatenated feature vector $F^{concat}$ is then flattened into a one-dimensional vector. The flattened feature vector is processed through a dense (fully connected) layer followed by a sigmoid activation function for binary classification.
3.3.5 Loss Function
We employ binary cross-entropy loss to train the classifier. This loss function measures the discrepancy between predicted probabilities and true labels for binary classification tasks (Equation 9).
$\text{Loss}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\cdot\log(\hat{y}_{i})+(1-y_{i})\cdot\log(1-\hat{y}_{i})\right]$ (9)
where $N$ is the number of samples, $y_{i}$ is the true label (0 or 1), and $\hat{y_{i}}$ is the predicted probability.
4 Experimental Result
4.1 Dataset
In this research, we utilized several datasets to effectively train and evaluate our models for both segmentation and classification tasks. For training and validating the TransUNet model, we used the ”Chest Xray Masks and Labels” dataset candemir2013lung ; jaeger2013automatic . This dataset contains 714 chest X-ray images, accompanied by their masks. Due to data limitation, our segmentation model was trained on 690 images and their corresponding masks, with 24 images reserved for validation purposes.
For classification task, we used two datasets. COVID-19 Image Data Collection provided by ”Cohen” cohen2020covid . This dataset comprises a total of 6,432 images, including three classes: Pneumonia, COVID-19, and Normal. The dataset is notably challenging due to its class imbalance and the complexity introduced by the three distinct classes. The distribution of images in the training set is as follows: 3,418 images of Pneumonia, 1,266 images of Normal, and 460 images of COVID-19. Approximately 20% of the images are allocated for testing.
Pediatric Pneumonia Chest X-ray Dataset provided by Kermany et al. kermany2018labeled . This dataset includes 5,856 images, with 5,232 images used for training and the remaining images reserved for testing. The dataset presents a significant challenge due to its class imbalance, with 3,883 images labeled as Pneumonia and 1,349 images as Normal. Additionally, the images in this dataset are from children, who often experience discomfort during X-ray procedures. This discomfort can impact the quality and consistency of the images, and the physiological differences between children and adults add an extra layer of complexity to the classification task. Both datasets contribute valuable and complementary challenges to our classification task, ensuring that our model is robust and capable of handling various real-world scenarios.
4.2 Experimental Setting
We implemented our proposed method using PyTorch version 1.8.1. In classification task, we utilized pre-trained ResNet-50 and ResNet-101 models, which were kept frozen during training to preserve their learned representations. Notably, the learnable parameters of our method amount to only 0.27 million. The model was trained for 30 epochs, and this process was repeated five times to ensure the robustness of the results. The average of these results was then reported.
The Adam optimizer was employed for training, with a learning rate set to $10^{-5}$. All input images were resized to 512 × 512 pixels, and the batch size was set to 64. Notably, no data augmentation techniques were applied; the results reported here were achieved without any such enhancements. All experiments were conducted on an NVIDIA RTX 4090 GPU.
Table 1: Performance on ”Cohen” dataset. Numbers in bold represent the best performance, while underlined values denote the second-best performance.
Models Acuracy Precision Recall F1-score
Densenet121* 87.8% 53.9% 71.0% 61.27%
Densenet169* 87.1% 32.3% 65.6% 43.28%
Densenet201* 88.4% 51.9% 79.0% 62.64%
Mobilenet_v2* 86.9% 33.4% 75.0% 46.21%
ResNet-50* 87.1% 38.4% 71.0% 49.84%
ResNet-101* 87.9% 33.5% 73.0% 45.92 %
Goodwinet al. (Ensemble learning) goodwin2020intra 89.4% 53.3% 80.0% 63.97%
Gadza et. al gazda2021self 84.9% 77.4% 90.6% 83.48%
Proposed method (ResNet-50 as backbone) 95.11% 95.81% 94.04% 94.92%
Proposed method (ResNet-101 as backbone) 94.18% 94.76% 92.79% 93.76%
Models marked with ”*” have results directly reported from goodwin2020intra .
Table 2: Performance on ”Kermany” dataset. Numbers in bold represent the best performance, while underlined values denote the second-best performance.
Models Acuracy Precision Recall F1-score
Yadav et al. (VGG16 as backbone) yadav2019deep 88.50% - - -
Ayan et al. (VGG16 as backbone) ayan2019diagnosis 87.98% 82.72% 85.90% 84.28%
Chattopadhyay et al. chattopadhyay2023exploring 81.7% - - 80.6%
Bhatt et al. (CNN) bhatt2023convolutional 85.58% 83.33% 96.15% 89.29%
Reshan et al. (MobileNet as backbone) reshan2023detection 0.85% 91.41% 95.28% 91.41%
Proposed method (ResNet-50 as backbone) 91.03% 91.54% 94.36% 92.93%
Proposed method (ResNet-101 as backbone) 92.79% 93.45% 95.13% 94.28%
Table 3: Effect of segmentation on ”Cohen” dataset
Backbones Results of proposed method Acuracy Precision Recall F1-score AUC
ResNet-50 on original images 91.23% 90.94% 85.60% 88.19% 89.41%
on predicted masks 95.11% 95.81% 94.04% 94.92% 95.07%
ReNet101 on original images 90.22% 88.16% 87.04% 87.6% 90.16%
on predicted masks 94.18% 94.76% 92.79% 93.76% 94.28%
4.3 Evaluation Metrics
To evaluate the effectiveness of our proposed approach, we use several performance metrics, including:
Accuracy: This metric reflects the ratio of correctly identified instances to the total number of instances. It is determined using Equation 10.
Accuracy=True Positives+True NegativesTotal Number of $\text{Accuracy}=\frac{\text{True Positives}+\text{True Negatives}}{\text{TotalNumber of Instance}}$ (10)
Accuracy offers a broad overview of the classifier’s performance but can be deceptive when dealing with imbalanced datasets.
Precision: This metric quantifies the ratio of correctly predicted positive instances to the total predicted positives. It is represented by Equation 11.
Precision=True PositivesTrue Positives+True $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{TrueNegatives}}$ (11)
Precision is especially valuable when the consequence of false positives is significant.
Recall: Also referred to as Sensitivity or True Positive Rate, Recall measures the ratio of correctly predicted positive instances to the total actual positives. It is expressed by Equation 12.
Recall=True PositivesTrue Positives+False $\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$ (12)
Recall is vital in situations where the cost of missing a positive instance (false negatives) is high, ensuring that most positive instances are detected.
F1 Score: The F1 Score represents the harmonic mean of Precision and Recall, offering a balance between these two metrics. It is particularly advantageous when handling imbalanced datasets. The F1 Score is determined using Equation 13.
[…]
This score provides a single metric that accounts for both false positives and false negatives, reflecting the classifier’s overall performance.
(a) Result of the proposed method on ResNet-50 with original images.
(b) Result of the proposed method on ResNet-50 with predicted masks.
(c) Result of the proposed method on ResNet-101 with original images.
(d) Result of the proposed method on ResNet-101 with predicted masks.
Figure 3: Confusion matrix of ”Cohen” dataset with different scenarios
4.4 Comparison with State-of-the-Art
We compared our proposed method with several state-of-the-art methods on both datasets. As shown in Table 1 and Table 2, our proposed method achieved the best results across all evaluation metrics, including accuracy, precision, recall, and F1 score, using both ResNet-50 and ResNet-101 backbones. This demonstrates the effectiveness of our approach in classifying pneumonia from chest X-ray images.
Our method leverages the strengths of pre-trained ResNet models to extract robust feature representations. By freezing the pre-trained layers, we focus the learning on the newly added layers, reducing the number of learnable parameters to only 0.27 million.
4.5 Ablation study
To evaluate the impact of the segmentation component on the performance of our classification model, we conducted an ablation study using the ”Cohen” dataset. This study compares the classification results obtained with and without the segmentation step, providing insights into the effectiveness of incorporating lung masks generated by the TransUNet model.
The ablation study involves evaluating the classification performance of our model with two different input scenarios: 1) Original Images: Classification is performed directly on the raw chest X-ray images from the ”Cohen” dataset. 2) Predicted Masks: Classification is performed on the chest X-ray images after segmenting the lung regions using the TransUNet model. The images used for classification are limited to the areas highlighted by the predicted lung masks.
The results of the ablation study are summarized in Table 3. The table displays classification metrics, including accuracy, precision, recall, F1-score, and AUC, for both ResNet-50 and ResNet-101 backbones under the two different input scenarios.
Table 4: The effect of using each key component on ”Cohen” dataset (on original images and without segmentation)
Backbones Baseline
Multi-scale
feature maps
Transformer Accuracy Precision Recall F1-score
ResNet-50 X 84.62% 75.68% 75.10% 75.39%
X X 87.73% 80.55% 79.62% 80.08%
X X 87.73% 80.63% 80.47% 80.55%
X X X 91.23% 90.94% 85.60% 88.19%
ResNet-101 X 83.93% 74.21% 73.92% 74.06%
X X 85.56% 78.21% 80.4% 79.29%
X X 85.71% 78.59% 81.11% 79.83%
X X X 90.22% 88.16% 87.04% 87.6%
The results clearly demonstrate the benefit of incorporating segmentation masks in the classification process. For both ResNet-50 and ResNet-101 backbones, the model achieves higher accuracy, precision, recall, F1-score, and AUC when trained on images with predicted lung masks compared to the raw images. Specifically, the accuracy improves by 2 percentage points for ResNet-50 and by 2.73 percentage points for ResNet-101. Similarly, the precision, recall, F1-score, and AUC all show substantial improvements.
These findings underscore the effectiveness of the segmentation component in isolating relevant features within the lung regions, which enhances the classification model’s ability to accurately diagnose pneumonia. By focusing on the segmented lung areas, the classification model benefits from reduced noise and more relevant information, leading to better overall performance.
In addition to the classification metrics presented in Table 3, we further analyze the performance of our model using confusion matrices under different scenarios, as depicted in Figure 3. The confusion matrices provide a detailed breakdown of the classification results, showing the true positives, true negatives, false positives, and false negatives for each class.
For the ResNet-50 backbone, the confusion matrix in Figure 3(a) shows the results on original images. The model correctly classifies 88 COVID-19 cases, 271 normal cases, and 816 pneumonia cases, with a notable number of misclassifications, particularly in the COVID-19 and pneumonia categories. When using predicted masks, as shown in Figure 3(b), the model’s performance improves significantly, correctly classifying 112 COVID-19 cases, 279 normal cases, and 834 pneumonia cases. The number of misclassifications decreases across all categories, highlighting the benefit of segmentation in isolating relevant features.
For the ResNet-101 backbone, the confusion matrix in Figure 3(c) displays the results on original images, where the model correctly classifies 94 COVID-19 cases, 278 normal cases, and 790 pneumonia cases. However, the misclassifications are more pronounced compared to ResNet-50, particularly in the pneumonia category. With the predicted masks, as shown in Figure 3(d), the performance improves, with correct classifications of 107 COVID-19 cases, 286 normal cases, and 820 pneumonia cases. This reduction in misclassifications further supports the effectiveness of incorporating segmentation masks.
These visual representations in the confusion matrices clearly demonstrate the improvement in classification performance when using the predicted masks. The consistent reduction in false positives and false negatives across both backbones underscores the robustness of the segmentation approach. By focusing on the lung regions and eliminating irrelevant background information, the segmentation component enhances the model’s ability to accurately diagnose pneumonia, resulting in better overall performance.
Furthermore, we investigate the contributions of key components within our model on the ”Cohen” dataset, specifically focusing on the impact of multi-scale feature maps and the transformer module. The results are presented in Table 4, highlighting the performance metrics, including accuracy, precision, recall, and F1-score, across different configurations of the ResNet-50 and ResNet-101 backbones. Each row of the table demonstrates how the incorporation of each component affects the model’s performance, providing a comprehensive view of their individual and combined effects.
The findings reveal a significant enhancement in model performance when both multi-scale feature maps and the transformer are employed alongside the baseline configuration. For instance, with the ResNet-50 backbone, the accuracy improves from 84.62% (baseline) to 91.23% when all components are utilized. Similarly, the ResNet-101 backbone exhibits a notable increase in accuracy from 83.93% to 90.22%. These results underscore the effectiveness of our proposed innovations, illustrating that the integration of multi-scale feature maps and transformer elements not only enhances overall accuracy but also boosts precision, recall, and F1-score, which are crucial for the reliability of classification tasks. This highlights the importance of these key components in achieving improved performance in deep learning models for image analysis.
5 Conclusion
This paper presents an innovative and efficient method for pneumonia detection utilizing a novel multi-scale transformer approach. By integrating lung segmentation using the TransUNet model with a specialized transformer module, our approach effectively isolates lung regions, thereby enhancing the performance of subsequent classification tasks. The proposed method demonstrates significant improvements in classification metrics, as evidenced by the ablation study on the ”Cohen” dataset. Both ResNet-50 and ResNet-101 backbones benefited from the segmentation masks, showing increased accuracy, precision, recall, and F1-score. These improvements underscore the effectiveness of our approach in focusing on relevant lung features while reducing noise from irrelevant regions.
The high accuracy rates of 92.79% on the ”Kermany” dataset and 95.11% on the ”Cohen” dataset confirm the robustness and reliability of our model. The reduction in the number of parameters compared to other state-of-the-art transformer models highlights our contribution to creating a more efficient yet powerful diagnostic tool suitable for deployment in resource-constrained environments. Our work paves the way for future research in several areas. Future work could explore further optimization of the transformer module to enhance performance and reduce computational complexity. Additionally, expanding the dataset to include a broader variety of pneumonia cases and other respiratory diseases could improve the model’s generalization.